OpenStudy (anonymous):

Put the following equation in slope-intercept form: y - 5 = (x - 3)

OpenStudy (anonymous):

Slope intercept form: y=mx+b y-5= x-3 get y alone on this side Can you do it?

OpenStudy (anonymous):

y=5x-3?

OpenStudy (anonymous):

Close! Dont forget that you have to add 5 to both x and -3. :o)